OpenStudy (anonymous):

Are lines and AB and CD parallel, perpendicular, or neither? A(-1,-4) B(2,11) C(1,1) D(4,10)

OpenStudy (anonymous):

@jollysailorbold

OpenStudy (anonymous):

I would say neither. @jollysailorbold

OpenStudy (anonymous):

do you want equations for it? or no?

OpenStudy (anonymous):

No. I had parallel first, then thought it was neither....idk ohwell lol

OpenStudy (anonymous):

haha, okay it's up to you.

OpenStudy (anonymous):

You have no clue what it is? what i should lean to

OpenStudy (anonymous):

they definately cross, but not at a 90 degree angle

OpenStudy (anonymous):

Perpendicular then!,,lol

OpenStudy (anonymous):

NOT at a 90 degree angle :)

OpenStudy (anonymous):

if the lines are perpendicular to each other, then the product of their slopes is -1 m1 * m2 = -1

OpenStudy (anonymous):

the slope of AB is 5, and the slope of CD is 9

OpenStudy (anonymous):

So whats that tell me?

OpenStudy (anonymous):

5 * 9 doesn't equal -1. so it's not perpendicular, or parellel.

OpenStudy (anonymous):

That makes much more sense now then! thats why i was stuck lol!

OpenStudy (anonymous):

:) glad i could help. haha

OpenStudy (anonymous):

Lol yeah totally! Can you help with a few more or no?

OpenStudy (anonymous):

um sure:)

OpenStudy (anonymous):

Okay:) having trouble uploading this picture lol

OpenStudy (anonymous):

just put in a link

OpenStudy (anonymous):

What you mean?

OpenStudy (anonymous):

you do online schooling yes? click on the picture you want, and just drag it to another tab, it should work:)

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

okay, so those are vertical angles, which mean that they are opposite from each other, and which also means that they equal the same amount. so 169 = 3x 169/3 = x which is 56 and 1/3 so x = 56 1/3 :) hope i helped

OpenStudy (anonymous):

Yeah makes sense.. but the answer just doesn't look right becuase thats what i had too! lol

OpenStudy (anonymous):

do you need it in a decimal form? 56 1/3 is the same as 56.3333333333333(repeating)

OpenStudy (anonymous):

Just doesn't look right for some reason..lol

OpenStudy (anonymous):

WAit! yeah your right.. nevermind:)

OpenStudy (anonymous):

solve for x.

OpenStudy (anonymous):

This ones confusing me.

OpenStudy (anonymous):